Please read these instructions before posting any event on Fermilab Indico

Indico will be unavailable due to server patching on Wed , June 19th, 7 a.m. to 7:30 a.m. Central Time

ND-GAr: HPgTPC+ECAL Weekly Meeting

US/Central
Disappearance Room (WH13SE) and Zoom

Disappearance Room (WH13SE) and Zoom

    • 10:00 AM 10:20 AM
      Roundtable Updates 20m
      Speakers: Alfons Weber (JG-Uni Mainz & Fermilab), Diego Gonzalez Diaz, Naseem Khan, Tanaz Mohayai (Fermilab)

      GOAT:

      Setting up second ALICE IROC.  Prepping to test it at 1 atm with argon/CO2 and a high dielectric gas to check voltage holdoff quickly.  If using SF6, need to collect it since it is a greenhouse gas.  Looking into alternatives. 

       

      GORG:

      Tanaz will move it to a cleanroom to open it up.  Maybe lab 6 or SciDet.

       

      TOAD:

      Naseem:  Met with Steve and hydraulic motor is working, but some discussion about hydraulic operation procedure.  Nearly ready for a pressure test (using existing procedure).  A source has been identified.  A class 3 55-Fe course has been identified, so don't need to borrow GOAT source.  Spellman HV part replaced to run on 115 V.  Naseem and Ioannis debugging TIP card.    Todd says gas has been delivered to DAB, and will request to move it to FTBF.    Have ordered 92/8 and pure Argon. 

      Alysia:  Was talking Joe yesterday about the wire chambers and beam instrumentation.  3 are in place, a 4th needs to be installed.  Joe said that they have gas (85%Argon/15% Iso) bottles.  Readout cables are in place, but need to install data acquisition box (they have some).  The 4 chambers also need a trigger, probably from scintillators that we need to have in beam. Will also need to run HV from the control room to the wire chambers.  No HV supplies in control room now (cannot use a "Berkeley cow" since need something with a current limiter).  Joe says they have lots of Bertan ones, and Jen thinks they had some nice remotely controllable ones for LAriat.    There seems to be a TOF box and a scintillator paddle upstream right now and maybe parts for a downstream TOF counter.  But will need to determine PMT voltages and check for light leaks etc.  Will also need to mock up some trigger logic in a NIM bin.  Eventually need to figure out how to combine these two data streams.   Also Joe showed us where the tclock signals come into the enclosure from the control room, so that we hopefully now know where to grab the timing info.

      Patrick:  Order for full 150-something boards is in. 

      Anezka:  No reco updates:

       

      Light Collection:

      Diego: See signals in SiPMs, but will make a new board.  4x4 SiPM tile and want to test them individually.  The TPC is up and running.  Using PMTs for primary scintillation right now.  Also see copious scintillation s2 with thick GEMs.  Started to implement the TPC in Geant4.  Also taking steps to increase optical gain of structures.  Trying to get new glass GEMs.

      Mirald:  For TPC optical simulation, where can latest version of ECAL geometry be found.  Francisco posted location of a geometry file in the chat.

       

      Software:

      Francisco - No major updates on CAFs from last week. But can now read the CAF files.    Input right now is a Genie GHEP file.